a sample of hydrogen has an inital twmprature of 50.0 celcius. when the temperature is lowered two -5.0°C the volume of hydrogen becomes 212 cm3 . what was the inital volume of the hydrogen?


1
Expert's answer
2021-06-11T01:57:51-0400

V1/T1 = V2/T2

V1/323 = 212/268

V1/323 = 0.7991

V1 = 0.7991 × 323

V1 = 255.51cm³